RC meeting and Run Plan, Wednesday Oct. 4, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)
After the RC meeting yesterday, Ellen went down into the hall to disconnect the RTDs that Gordon had asked for. Beam was restored at 6:30pm, and ran at 45uA until Owl shift. Issues were raised with the RC about poor statistics in the 50k replays for the first 2 runs, and a supposed second spot on the beam-on-target plots (this has been resolved, and was due to misunderstanding the plots). In addition, the beam dropped for a BLM trip at 2:15am; when the TO tried to move to No Target at the behest of MCC, we found that IOCHA26 had crashed. It was rebooted by 4:20am, and Owl shift continued to run 45uA. That has been maintained during Day shift, minus 20 minutes to restart the SBS GEM HV, and 10 minutes to again reboot IOCHA26 when the oven flow and temperature plots flatlined.
- No major issues with the accelerator today.
- This morning, MCC declared that they will have some available access time in the next 2 weeks.
- Beam studies will take place during Day shift tomorrow (10/5)
- There will be a maintenance day during Day shift on Wednesday 11/10
If you want to get into the hall, please communicate with the RC to get this organized!
Data collected :
- Accumulated corrected charge collected: 1.216C across the three shifts since yesterday's Day shift.
Hall A Subsystems Status
- IOCHA26 has been restarted twice today; once to fix the target ladder not responding, and once to fix a readback issue with the oven temperature and flow.
- The SBS GEM HV was restarted a couple of times this morning to address trips. Holly has increased the GEM HV trip points slightly to try to prevent this happening again.
Hall access requests
- If anyone wants to make an access, let the RC know!
- All shifts have been filled until 10/12.
- Please check when your shifts are!
Please sign up for any open shifts!
Short Term Run plan
- Standard production running at 45uA, with NMRs every 3 hours.
==> List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)